The membrane that surrounds a certain type of living cell has a surface area of 4.70 x 10-9 m2 and a thickness of 9.60 x 10-9 m. Assume that the membrane behaves like a parallel plate capacitor and has a dielectric constant of 5.40 (a) The potential on the outer surface of the membrane is +62.4 mv greater than that on the inside surface. How much charge resides on the outer surface? 1.11e-12X (b) If the charge in part (a》 is due to K+ ions (charge +e), how many such ions are present on the outer surface? 6.94e6 ions
Answer #1

(a)

we know

C = k*\varepsilono*A/d

C = 5.40*8.85*10-12*4.70*10-9/9.60*10-9

C = 2.33*10-11 F

q = V*C

q = 62.4*10-3*2.33*10-11

q = 14.59*10-13 C

(b)

number of ions n = q/e = 14.59*10-13/1.6*10-19

n = 9.12*106